Sub 25
Ques. 14

Simplify:
(82/3 - 81/3)3
note: 2/3, 1/3, 3 are exponents
1
Expert's answer
2012-05-11T08:10:58-0400
(8^2/3 - 8^1/3)^3=(4-2)^3=8